How do you insert grouping symbols in this equation to make it true? 8+ 25 x 2 divided by 6=11 ?

Answer by jim_thompson5910(35256)

Place parenthesis around 8 and 25 to get


(8+25)x2/6 = 11
33x2/6 = 11
66/6 = 11
11 = 11


If you need more help, email me at jim_thompson5910@hotmail.com

Also, feel free to check out my tutoring website

Jim